Grade 12th passMechanics

An engine is working at a constant power draws a load of mass m against a resistance R. Find maximum speed of load and time to attain half this speed..... Please provide complete solution. .........

Pawan joshi
7 Years agoGrade 12th pass

1 Answer

Piyush Kumar Maurya
7 Years ago
Let maximum speed of load be v.
Power P= rv.
v = P/r
Let X be the distance traversed from initial time to time when it acquires maximum speed
m *dv/dt = r = P/v
m*vdv = Pdt
Integrating from 0 to v/2 i.e. P/2r on LHS and 
0 to t on RHS
1/2 *m *P/ 4r= PT
t = Pm/8r2